OpenStudy (anonymous):

media service charges $40 for a phone and $20/month for it economy plan .the equation c=20t=40 describes the total cost c,of operating a media service phone for t month

OpenStudy (anonymous):

What exactly is the question?

OpenStudy (anonymous):

the total cost for 4 month of service is ?

OpenStudy (anonymous):

$120

OpenStudy (anonymous):

if a customer has only $90 available how many months of service can she receive

OpenStudy (anonymous):

2 months, 20*2=40. 40+40=80.

OpenStudy (anonymous):

After the 120 do i need to put 120t

OpenStudy (anonymous):

No, cus then t=1. not the case.... Are you plugging it into the equation?

OpenStudy (anonymous):

Your equation is faulty. if the total cost is c, then the equation should be c = 40 + 20t.